OpenStudy (anonymous):

Two sides of a triangle have lengths 9 m and 15 m. The angle between them is increasing at a rate of 2°/min. How fast is the length of the third side increasing when the angle between the sides of fixed length is 60°? (Round your answer to three decimal places.)

OpenStudy (anonymous):

guess you need the law of cosines for this relationship do you know it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

9^2+15^2-2(9)(15)cos

OpenStudy (anonymous):

if you call the side you are interested n \(x\) you have \[x^2=15^2+9^2-2\times 15\times 9\cos(\theta)\]

OpenStudy (anonymous):

ok then take the derivative wrt time and get \[2xx'=270\sin(\theta)\theta'\]

OpenStudy (anonymous):

you are told \(\theta'=2\) and that \(\theta=60\) plug them in and solve

OpenStudy (anonymous):

actually there is something seriously wrong with this problem, but i won't bother you about it i will just add it to my "bad math" pile

OpenStudy (anonymous):

i would love to see what answer you get i can guarantee that it will not be correct

OpenStudy (anonymous):

270 sqrt(3)

OpenStudy (anonymous):

i am wondering who wrote this problem

OpenStudy (anonymous):

it has a big flaw in it, or else we need to do something much more complicated

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you happen to know what the right answer is supposed to be?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i got 0.416

OpenStudy (anonymous):

ok well lets pretend that there is no problem here, i will get back to that you have \[2xx'=270\sin(\theta)\theta'\] which becomes \[2xx'=270\times \frac{\sqrt{3}}{2}\times 2\] but you still need \(x\) in order to solve for \(x'\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

solve for \(x\) here \[x^2=15^2+9^2-2\times 15\times 9\cos(60)\]

OpenStudy (anonymous):

i get \(x^2=171\) making \(x=\sqrt{171}\) or \(3\sqrt{19}\) if you prefer

OpenStudy (anonymous):

yes, i got that too

OpenStudy (anonymous):

so maybe the answer is supposed to be \[x'=\frac{270\sqrt{3}}{3\sqrt{19}}\] but i am sure that this is all nonsense for one simple reason

OpenStudy (anonymous):

sine and cosine are functions as function they are functions of numbers, not angles as functions of numbers they correspond to angles, only if the angles are measured in radians, not degrees as a function of degrees, the derivative of cosine is NOT sine so this is all junk

OpenStudy (anonymous):

2= pi/90

OpenStudy (anonymous):

convert it into radians.

OpenStudy (anonymous):

ok then it should work of course you get a completely different answer

OpenStudy (anonymous):

it is a really really bad idea to use calculus and degrees in the same problem they do not match up

OpenStudy (anonymous):

so what is the answer?

OpenStudy (anonymous):

i keep getting 0.416

OpenStudy (anonymous):

replace the \(2\) here by \(\frac{\pi}{90}\) as you said